Operations Report – March 4th

Crew 293 Operations Report 04-03-2024


Name of person filing report: Leo Tokaryev

Non-nominal systems: Science Dome and its tunnel

Notes on non-nominal systems: we found some water leaks on the Science Dome this morning.

Tunnel leading to the Science Dome is still not repaired. Today the wind was calm, but we didn’t have time to think about the problem because of the EVA. We’ll try to figure out a solution for the tunnel and keep you updated.


Spirit rover used: No

Hours: (before EVA)

Beginning charge: (Before EVA)

Ending charge: (On return from EVA, before recharging)

Currently charging:

Opportunity rover used: Yes

Hours: 165.1

Beginning charge:100

Ending charge:60

Currently charging:No

Curiosity rover used: Non

Hours: 269.7

Beginning charge:100

Ending charge:56

Currently charging:No

Perseverance rover used: No


Beginning charge:

Ending charge:

Currently charging:

General notes on rovers:none

Summary of Hab operations: put notes here

WATER USE: 15.6 gallons

Water (static tank): 270.5 gallons

Static tank pipe heater (on or off): ON

Static tank heater (On or off): ON

Toilet tank emptied: no

Summary of internet:

Summary of suits and radios:

Summary of GreenHab operations: put notes here

WATER USE: 8 gallons

Heater: On

Supplemental light: 10pm-2am

Harvest: 700g tomatoes, 5g basil

Summary of ScienceDome operations: None

Dual split: Heat or AC, On

Summary of RAM operations: (Tools used, work done):

Summary of any observatory issues: none

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support: we have left 3 trash bags in the rear airlock, could you please take care of it?

Copyright © The Mars Society. All rights reserved. | Main Site